MASM Assembly Language for x86 Processors:

Given the following array declaration: .data matrix DWORD 50 DUP(10 DUP(?)) If matrix[0][0] is the 0th sequentially stored BYTE in memory, which sequentially stored BYTE is the first byte corresponding to matrix[14][2]? (in decimal)

Please expplain how you got the answer, so I can answer other questions like this! Thank you
Answer #1

This is a 50 x 10 matrix, with each row taking 4 * 10 bytes.

Therefore to figure out the address of given row, add 14 * 4 * 10 = 560.
Next consider given column, add 2 * 4 = 8.


Ans: (14 * 4 * 10) + (2 * 4) = 568 bytes